OpenStudy (twalt13):

can someone help me ? The position of an object at time t is given by s(t) = -9 - 3t. Find the instantaneous velocity at t = 8 by finding the derivative

OpenStudy (ivancsc1996):

So you only need to derivate the equation:\[\frac{ ds(t) }{ dt }=v(t)=\frac{ d }{ dt }(-9-3t)=-3\]This means that the velocity is always three.

OpenStudy (ivancsc1996):

Sorry -3

OpenStudy (twalt13):

thanks